OpenStudy (haleyelizabeth2017):

Find the maximum value of the function for the polygonal convex set determined by the given system of inequalities.

OpenStudy (haleyelizabeth2017):

\(3x-2y\ge 0\) \(y\ge 0\) \(3x+2y\le 24\) \(f(x,y)=7y-3x\) Do I solve this one the same way as the last?

OpenStudy (haleyelizabeth2017):

...except find the maximum instead of minimum?

OpenStudy (phi):

yes

OpenStudy (haleyelizabeth2017):

Awesomeness...

OpenStudy (haleyelizabeth2017):

So...f(0,0)=7(0)-3(0) f(0,0)=0 f(4,6)=7(4)-3(6) f(4,6)=28-18 f(4,6)=10 f(8,0)=7(8)-3(0) f(8,0)=56 so the maximum is 56...

OpenStudy (phi):

isn't f(x,y)= 7y -3x ?

OpenStudy (haleyelizabeth2017):

shoo...yes -_-

OpenStudy (haleyelizabeth2017):

good catch...

OpenStudy (haleyelizabeth2017):

..so it's really 30..

OpenStudy (haleyelizabeth2017):

time to tackle my next assignment!